# On the Resilience of Multi-Agent Systems with Malicious Agents

### v2

@@ Acknowledgments
The paper is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund). The paper is also based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410.
Appendix A More Results
A.1 GPT-4o Results
(a) The three multi-agent system architectures.
(b) The four downstream tasks.
Figure 9: Performance drops of the six multi-agent systems on four selected downstream tasks.
To ensure a fair comparison with GPT-3.5 results, both AutoTransform and AutoInject use GPT-3.5, maintaining consistency with previous settings. Our conclusions remain valid for GPT-4o: (1) While performance improves across all structures, the “Hierarchical” structure demonstrates the highest resilience against malicious agents. (2) More rigorous tasks, such as code generation and solving math problems, experience greater performance declines. (3) We also observe a performance increase when using AutoInject across the three systems—MAD, Camel, and AgentVerse.
A.2 Quantitative Results
@@
GPT-3.5 63.10 63.70 67.40
          w/ AutoTransform 39.95 50.32 52.15
          w/ AutoInject 33.80 32.00 51.50
GPT-4o 82.60 77.75 84.76
          w/ AutoTransform 27.91 70.73 76.83
          w/ AutoInject 38.69 63.42 78.35
@@
Semantic 26.80 40.90 29.30 34.80 75.70 49.40 39.15
Table 5: Performance of our defense methods, the “Challenger” and “Inspector.” The percentages in brackets show the proportion of recovered performance loss calculated by $(X-A)/(V-A)\times 100\%$, where $A$ is the performance against AutoInject, $V$ is the vanilla performance, and $X$ is the performance with a specific defense method.
Setting Vanilla AutoInject Challenger Inspector
Self-collab 76.22 40.85 71.95 (87.93%) 67.68 (75.86%)
Camel 62.20 29.27 40.24 (33.33%) 44.16 (45.21%)
Appendix B Prompt Details
All six multi-agent collaboration systems selected in this study support only some of the downstream tasks in their original design. Therefore, we extend three scalable systems—Camel, MAD, and AgentVerse—to adapt to all four downstream tasks. The first three systems provide a high-level, non-task-oriented design for task division, while the other three systems are deeply intertwined with code generation tasks. Using Camel as an example of adapting systems to other tasks: For translation and math, we improve system performance by adding “step by step” instructions in prompts. For instance, in translation, it correctly interprets “{CJK}UTF8gkai拉下水 (pull into water)” to its correct meaning of “engaging in wrongdoing” in Chinese. In math, a single agent calculates “Average Speed$=(1+3)/2=1m/s$,” whereas Camel’s multi-agent system correctly computes “average speed$=(1+3)/2=2m/s$.” The detailed instructions likely reduce the occurrence of “seemingly” correct answers and increase accuracy in these specific cases.
B.1 Multi-Agent Systems on Different Tasks
B.1.1 Camel
@@ 2. An economist who is good at …
B.2 Single Agent on Different Tasks
@@ B.6 Defense Method: Inspector
Appendix C Limitations
There are several limitations in this study. First, due to budget constraints, we explore only GPT-3.5 and GPT-4o. Since our primary goal is to fairly evaluate different multi-agent systems’ resilience against malicious agents, we believe the results would not significantly differ from other models. The second limitation is the selection of multi-agent systems and downstream tasks, which cannot be comprehensive. We mitigate this by selecting representative systems from three well-established human collaboration modes Yang & Zhang (2019); Alexy (2022); Mihm et al. (2010) and using four commonly-used datasets for benchmarking the abilities of multi-agent systems Liang et al. (2024); Chen et al. (2021). The final limitation concerns the analysis, where latent variables affecting system resilience might be unidentified. To minimize this risk, we examine system architectures, downstream tasks, error rates, error types, agent roles, and the number of agents’ communications. To the best of our knowledge, no additional factors influencing system resilience are found.
Appendix D Ethics Statements and Broader Impacts
The two error introduction methods developed in this study, AutoTransform and AutoInject, could potentially pollute benign agents and result in negative social impacts. To mitigate this risk, we have proposed effective defense mechanisms against them. We would like to emphasize that the goal of proposing these methodologies is to study and improve the behavior of LLM-based multi-agent systems. We strongly oppose any malicious use of these methods to achieve negative ends.

### v3

@@ Acknowledgments
The paper is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410. The paper is also supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).
Appendix A Error Type Analysis
Category Name Description Count
Logical Errors Errors in logical operations, such as incorrect operators or inverted logic. 12
Indexing and Range Errors Issues with boundary conditions or off-by-one indexing. 23
Mathematical Errors Errors in calculations or numerical processing. 20
Output and Formatting Issues with producing or formatting expected output. 9
Initialization Errors Problems with starting values or incorrect initialization. 4
Infinite Loops Errors causing unintended infinite execution loops. 6
Runtime Invocation Issues Errors in function calls or runtime handling. 6
Table 4: Statistics of 80 errors injected by AutoInject in code generation.
We analyze the distribution of error types generated by AutoInject in code generation. The errors span across seven distinct categories, as detailed in Table 4, ensuring diversity in the types of faults injected and reducing the bias of any single category dominating the results. By incorporating a diverse range of errors and generating them at scale, AutoInject effectively captures the broad spectrum of fault types, mitigating the risk that specific critical cases—like infinite loops—are overlooked. This approach ensures that the reported error metrics, while simple, remain robust and representative of diverse error scenarios.
Appendix B Quantitative Results
@@
GPT-3.5 55.62 54.37 53.00
        w/ AutoTransform 38.24 43.93 46.57
        w/ AutoInject 38.27 40.25 48.12
GPT-4o 67.18 67.52 67.79
        w/ AutoTransform 30.08 56.92 60.83
        w/ AutoInject 44.14 60.51 64.01
@@
Syntactic 29.30 75.60 42.70 28.70 67.10 43.30 47.78
Appendix C Prompt Details
All six multi-agent collaboration systems selected in this study support only some of the downstream tasks in their original design. Therefore, we extend four scalable systems—MetaGPT, Camel, MAD, and AgentVerse—to adapt to all four downstream tasks. These systems provide a high-level, non-task-oriented design for task division, while the other two systems, namely Self-collab and SPP, are deeply intertwined with code generation tasks. Using Camel as an example of adapting systems to other tasks: For translation and math, we improve system performance by adding “step by step” instructions in prompts. For instance, in translation, it correctly interprets “{CJK}UTF8gkai拉下水 (pull into water)” to its correct meaning of “engaging in wrongdoing” in Chinese. In math, a single agent calculates “Average Speed$=(1+3)/2=1m/s$,” whereas Camel’s multi-agent system correctly computes “average speed$=(1+3)/2=2m/s$.” The detailed instructions likely reduce the occurrence of “seemingly” correct answers and increase accuracy in these specific cases.
C.1 Multi-Agent Systems on Different Tasks
C.1.1 MetaGPT
Prompt Template for MetaGPT
Engineer You are an expert in the field of $<$SUBJECT$>$, your goal is $<$GOAL$>$.
ATTENTION: Use ‘##’ to SPLIT SECTIONS, not ‘#’. Output format carefully referenced “Format example.”
# Context
## Design $<$DESIGN$>$
## Task $<$TASK$>$
## Legacy Results $<$LEGACY_RESULTS$>$
## Evaluation results $<$EVALUATION$>$
# Format example
## Deduction process and reasons (The reason for your answer)
## Answer (Your answer without further description, follow the format given in the task section)
# Instruction: Based on the context, follow “Format example,” write your answer below:
Reviewer You are an expert in the field of $<$SUBJECT$>$, your goal is $<$GOAL$>$
ATTENTION: Use ‘##’ to SPLIT SECTIONS, not ‘#’. Output format carefully referenced “Format example.”
# Context
## Design $<$DESIGN$>$
## Task $<$TASK$>$
## Legacy Results $<$LEGACY_RESULTS$>$
# Format example 1
## Review: 1. No, we should fix the logic in part … 2. … 3. No, there is some error in … 4. …
## Actions: 1. Fix the logic: The_fixed_solution 2. Revise the error: Sample_revised_version
## Review Result: LBTM
# Format example 2
## Review: 1. Yes. 2. Yes. 3. Yes. 4. Yes.
## Actions: Pass
## Review Result: LGTM
# Instruction: Based on the actual situation, follow one of the “Format example.” Return only 1 result for review.
## Review: Ordered List. Based on the “result to be Reviewed,” provide key, clear, concise, and specific answer. If any answer is no, explain how to fix it step by step.
1. Is the result implemented as per the requirements? If not, how to achieve it? Analyze it step by step.
2. Is the result logic completely correct? If there are errors, please indicate how to correct them.
3. Does the existing result contain any missing on edge cases?
4. Are all calculation correct? If there is no calculation, please indicate how to achieve it step by step.
5. Have the answer contain any subtle errors?
6. Are the Design being realized correctly?
## Review Result: str. If the result doesn’t have any errors, we don’t need to rewrite it, so answer LGTM and stop. ONLY ANSWER LGTM/LBTM.
# Instruction: Based on the context, follow “Format example,” write your answer below:
C.1.2 Camel
@@ 2. An economist who is good at …
C.2 Single Agent on Different Tasks
@@ C.6 Inspector
Appendix D Limitations
There are several limitations in this study. First, due to budget constraints, we explore only GPT-3.5 and GPT-4o. Since our primary goal is to fairly evaluate different multi-agent systems’ resilience against faulty agents, we believe the results would not greatly differ from other models. The second limitation is the selection of multi-agent systems and downstream tasks, which cannot be comprehensive. We mitigate this by selecting representative systems from three well-established human collaboration modes (Yang & Zhang, 2019; Alexy, 2022; Mihm et al., 2010) and using four commonly-used datasets for benchmarking the abilities of multi-agent systems (Liang et al., 2024; Chen et al., 2021). The final limitation concerns the analysis, where latent variables affecting system resilience might be unidentified. To minimize this risk, we examine system architectures, downstream tasks, error rates, error types, agent roles, and the number of agents’ communications. To the best of our knowledge, no additional factors influencing system resilience are found.
Appendix E Ethics Statements and Broader Impacts
The two error-introducing methods developed in this study, AutoTransform and AutoInject, could potentially pollute benign agents and result in negative social impacts. To mitigate this risk, we have proposed effective defense mechanisms, the Challenger and the Inspector, against them. We would like to emphasize that the goal of proposing these methodologies is to study and improve the behavior of LLM-based multi-agent collaboration. We strongly oppose any malicious use of these methods to achieve negative ends.